# Transferring spectroscopic stellar labels to 217 million Gaia DR3 XP stars with SHBoost

### v1

@@ C. Jordi
(Received July 9, 2024; accepted …)
@@ 1 Introduction
Recently, the scope of the Gaia data has been enlarged by the inclusion of more than 219 million low-resolution XP spectra (Carrasco et al. 2021; De Angeli et al. 2023; Montegriffo et al. 2023) in the third Gaia data release, DR3 (Gaia Collaboration et al. 2023c). For Gaia DR3, the mean XP spectra have primarily been used to infer much more precise stellar parameters like $T_{\rm eff},\log g,d,A_{0},$ and [M/H] (Andrae et al. 2023a; Fouesneau et al. 2023). Various groups have already shown that this new dataset can potentially be used for a variety of Galactic archaeology purposes: to estimate rough stellar $\alpha$-element abundances (Gavel et al. 2021; Hattori 2024; Laroche & Speagle 2024; Li et al. 2024), to efficiently find extremely metal-poor stars (Witten et al. 2022; Xylakis-Dornbusch et al. 2022, 2024; Lucey et al. 2023; Yao et al. 2024), to classify white dwarfs (Echeverry et al. 2022; Gaia Collaboration et al. 2023b), to study hydrogen emission and absorption lines (Weiler et al. 2023; Carrasco et al. 2024), or to improve the accuracy of stellar parameters determined from Gaia RVS spectra (Guiglion et al. 2024).
@@ 2.1 Spectroscopic labels
2.2 Training, validation, and test datasets
@@
ML algorithms (especially neural networks) are often susceptible to the varying quality of the training data. To test the impact of this, we have performed tests using different quality cuts for the input data, finding no significant changes. We therefore chose to use the full set of columns for each of the training data sets.
For each label (extinction $A_{V}$, effective temperature $\log T_{\rm eff}$, surface gravity $\log g$, metallicity [M/H], and stellar mass $M$), we apply a set of criteria to clean the training data from poorly determined or unreliable labels. These criteria are detailed in Table 1.
@@ 3.2 Typical accuracy and precision
Figure 3: Uncertainty distributions. Top row: Logarithmic histograms of the uncertainties for the training (yellow), test (green), and full Gaia DR3 XP (blue) samples. The median uncertainties are indicated in each panel. Second row: Uncertainties as a function of Gaia $G$ magnitude. Bottom row: Uncertainties as a function of distance.
@@
The typical precisions achieved for each label in the full Gaia DR3 XP sample are presented in Fig. 3. In particular, we achieve median precisions of 0.20 mag in $A_{V}$, 0.014 dex in $\log T_{\rm eff}$, 0.20 mag in $\log g$, 0.18 dex in [M/H], and 12% in mass. These uncertainties are highly heteroscedastic: for example, more than 50 million stars (using the 25th percentile of the label uncertainty distributions) have uncertainties smaller than 0.11 mag in $A_{V}$, 0.010 dex in $\log T_{\rm eff}$, 0.13 dex in $\log g$, 0.14 dex in [M/H], or 8% in mass, respectively. As expected, the uncertainty distributions for the full dataset (top row of Fig. 3) are broader than for the training and test data (since the bulk of the sample is fainter than most of the training set), but the typical uncertainties are still smaller than the ones obtained from pure astro-photometric isochrone fitting (as performed in e.g. Anders et al. 2022; see Appendix A.2). As expected, the reported uncertainties increase with magnitude for most of the labels (except for stellar mass, which is also expected). The uncertainties also increase with distance, as expected.
@@ 5.3 Extinction maps
Figure 10: Differential $A_{V}$ extinction maps for the Orion region (30 deg x 30 deg, $d<1.5$ kpc; left) in four bins of distance, as indicated in each panel. Some artifacts from the Gaia scanning law are visible in these maps, but in general the distance-sliced $A_{V}$ maps are complete down to $A_{V}\gtrsim 6$ mag, have a high resolution, and and correspond very well to existing three-dimensional dust maps.
@@ 6 Conclusions and lessons for future work
Our new SHBoost catalogue comes with reliable uncertainties estimated via xgboost-distributions. We achieve median (best 25% percentile in brackets) precisions of 0.20 (0.11) mag in line-of-sight extinction $A_{V}$, 0.014 (0.010) dex in effective temperature $\log T_{\rm eff}$, 0.20 (0.13) mag in surface gravity $\log g$, 0.18 (0.14) dex in metallicty [M/H], and 12% (8%) in mass. The validation of the individual uncertainty estimates shows that they are correctly estimated. In addition, our xgboost models can be interpreted by using SHAP values (see Sect. 3.3 and Fig. 4).
@@
The whole inference pipeline (building the training set, training of the xgboost regression of stellar parameters, post-processing) can be run in 2 days on a 48-core machine at the Leibniz-Institut für Astrophysik Potsdam. This is a significant improvement in computational speed compared to the latest StarHorse EDR3 run, which took roughly three weeks to complete on a similar machine. We emphasize, however, that the need for classical Bayesian inference (at least for the construction of lare robust training sets) will persist in the future.
@@ Appendix B Data model and usage examples
4 xgb_teff K effective temperature, xgboost point estimate
5 xgb_teff_mean K effective temperature, xgboost-distribution mean value
6 xgb_teff_std K effective temperature, xgboost-distribution standard deviation

### v2

@@ C. Jordi
(Received September 18, 2024; accepted …)
@@ 1 Introduction
Recently, the scope of the Gaia data has been enlarged by the inclusion of more than 219 million low-resolution XP spectra (Carrasco et al. 2021; De Angeli et al. 2023; Montegriffo et al. 2023) in the third Gaia data release, DR3 (Gaia Collaboration et al. 2023c). For Gaia DR3, the mean XP spectra have primarily been used to infer much more precise stellar parameters like $T_{\rm eff},\log g,d,A_{0},$ and [M/H] (Andrae et al. 2023a; Fouesneau et al. 2023). Various groups have already shown that this new dataset can potentially be used for a variety of Galactic archaeology purposes: to estimate rough stellar $\alpha$-element abundances (Gavel et al. 2021; Hattori 2024; Laroche & Speagle 2024; Li et al. 2024), to efficiently find extremely metal-poor stars (Witten et al. 2022; Xylakis-Dornbusch et al. 2022, 2024; Lucey et al. 2023; Yao et al. 2024), to classify white dwarfs (Echeverry et al. 2022; Gaia Collaboration et al. 2023b), to study hydrogen emission and absorption lines (Weiler et al. 2023), or to improve the accuracy of stellar parameters determined from Gaia RVS spectra (Guiglion et al. 2024).
@@ 2.1 Spectroscopic labels
2.2 Training and test datasets
@@
ML algorithms (especially neural networks) are often susceptible to the varying quality of the training data. To test the impact of this, we have performed tests using different quality cuts for the input data (i.e. varying the maximum uncertainty for each parameter; see second line in Table 1), finding no significant improvements. We therefore chose to use the full set of columns for each of the training data sets.
For each label (extinction $A_{V}$, effective temperature $\log T_{\rm eff}$, surface gravity $\log g$, metallicity [M/H], and stellar mass $M$), we apply a set of criteria to clean the training data from poorly determined or unreliable labels, as detailed in Table 1. For each label, 80% of the data are used during the training phase (see next section), while 20% of the data are reserved for estimating the overall performance (test dataset; see Fig. 2).
@@ 3.2 Typical accuracy and precision
[…]
@@
The typical precisions achieved for each label in the full Gaia DR3 XP sample are presented in Fig. 3. In particular, we achieve median precisions of 0.20 mag in $A_{V}$, 0.014 dex in $\log T_{\rm eff}$ (or 174 K in $T_{\rm eff}$), 0.20 mag in $\log g$, 0.18 dex in [M/H], and 12% in mass. These uncertainties are highly heteroscedastic: for example, more than 50 million stars (using the 25th percentile of the label uncertainty distributions) have uncertainties smaller than 0.11 mag in $A_{V}$, 0.010 dex in $\log T_{\rm eff}$, 0.13 dex in $\log g$, 0.14 dex in [M/H], or 8% in mass, respectively. As expected, the uncertainty distributions for the full dataset (top row of Fig. 3) are broader than for the training and test data (since the bulk of the sample is fainter than most of the training set), but the typical uncertainties are still smaller than the ones obtained from pure astro-photometric isochrone fitting (as performed in e.g. Anders et al. 2022; see Appendix A.2). As expected, the reported uncertainties increase with magnitude for most of the labels (except for stellar mass, which is also expected). The uncertainties also increase with distance, as expected.
@@ 5.3 Extinction maps
Figure 10: Differential $\Delta A_{V}$ extinction maps for the Orion region (30 deg x 30 deg, $d<1.5$ kpc; left) in four bins of distance, as indicated in each panel. Some artifacts from the Gaia scanning law are visible in these maps, but in general the distance-sliced $A_{V}$ maps are complete down to $A_{V}\gtrsim 6$ mag, have a high resolution, and and correspond very well to existing three-dimensional dust maps.
@@ 6 Conclusions and lessons for future work
Our new SHBoost catalogue comes with reliable uncertainties estimated via xgboost-distributions. We achieve median (best 25% percentile in brackets) precisions of 0.20 (0.11) mag in line-of-sight extinction $A_{V}$, 0.014 (0.010) dex in effective temperature $\log T_{\rm eff}$ (or 174 [119] K in $T_{\rm eff}$), 0.20 (0.13) mag in surface gravity $\log g$, 0.18 (0.14) dex in metallicty [M/H], and 12% (8%) in mass. The validation of the individual uncertainty estimates shows that they are correctly estimated. In addition, our xgboost models can be interpreted by using SHAP values (see Sect. 3.3 and Fig. 4).
@@
The whole inference pipeline (building the training set, training of the xgboost regression of stellar parameters, post-processing) could be run in 2 days on a 48-core machine at the Leibniz-Institut für Astrophysik Potsdam. This is a significant improvement in computational speed compared to the latest StarHorse EDR3 run, which took roughly three weeks to complete on a similar machine. We emphasize, however, that the need for classical Bayesian inference (at least for the construction of large robust training sets) will persist in the future.
@@ Appendix B Data model and usage examples
4 xgb_logteff [K] effective temperature, xgboost point estimate
5 xgb_logteff_mean [K] effective temperature, xgboost-distribution mean value
6 xgb_logteff_std [K] effective temperature, xgboost-distribution standard deviation